# Distributed memory Tensor-train crossT. Shi, D. Hayes, and J. Qiu

### v1

@@ 6 Conclusions
We introduce a new distributed memory subtensor parallel algorithm for constructing a TT cross approximation for a given tensor. This construction allows for local computations with little communication requirements to obtain TT cores at a global level, while maintaining the interpolation property of a standard TT cross algorithm. This algorithm relies on the efficient update method for pivot selection using the material of Section 3.1.2, which ensures the selection of the optimal global pivot for each dimension. Furthermore, we utilize multiple process grids of Section 3.2.3 combined with an alternate recursive update formula Section 3.2.2 to construct all TT core simultaneously, with each core constructed in a distributed memory framework. Our presented numerical results demonstrate results ranging from good to optimal scaling for both 3d and 6d Hilbert tensors as well as real world 4d and 6d Maxwellian datasets.

### v2

@@ 6 Conclusions
We introduce a new distributed memory subtensor parallel algorithm for constructing a TT cross approximation for a given tensor. This construction allows for local computations with little communication requirements to obtain TT cores at a global level, while maintaining the interpolation property of a standard TT cross algorithm. This algorithm relies on the efficient update method for pivot selection using the material of Section 3.1.2, which ensures the selection of the optimal global pivot for each dimension. Furthermore, we utilize multiple process grids of Section 3.2.3 combined with an alternate recursive update formula Section 3.2.2 to construct all TT core simultaneously, with each core constructed in a distributed memory framework. Our presented numerical results demonstrate results ranging from good to optimal scaling for both 3d and 6d Hilbert tensors as well as real world 4d and 6d Maxwellian datasets. Furthermore, in comparisons with a dimension parallel algorithm for index selection, as well as publicly available package torchTT, we observe significantly lower run time in all test cases.